A solution contains 57.0 g octane, C_8 H_18 (M = 1
Answer #1

1. molality of mixer(m) = n/W in kg

   n = no of mole of solute(octane) = w/Mwt = 57/114 = 0.5 mole

W = mass of benzene(solvent) in kg = 0.148 kg

m = 0.5/0.148 = 3.38 molal

answer: d

2. molarity of solution = n/V

n = no of mole of solute(octane) = w/Mwt = 57/114 = 0.5 mole

v = volume of solution in L

mass of solution = 57+148 = 205 g

volume of solution = mass/density = 205/0.83 = 247 ml

molarity(M) = 0.5/0.247 = 2.02 M

answer: 2.02 M

3. DTf = i*kf*m

(T0 - Ts) = i*kf*m

i = vanthoff factor = 1

kf of benzene = 5.12 c/molal

m = molality of mixer = n/W in kg = 3.38 m

T0 = freezing point of solvent(C6H6) = 5.5 c

Ts = freezing point of solution = ?

(5.5-x) = 1*5.12*3.38

x = -11.8 C


freezing point of solution = -11.8 C